# Improving Diffusion Inverse Problem Solving with Decoupled Noise Annealing

### v2

@@ 4.2 Main Results
In addition, our method can generate diversified samples when the measurements contain less information. To illustrate this, we selected two tasks where the posterior distributions may have multiple modes: (1) super-resolution by a factor of 16, and (2) box inpainting with a box size of 192$\times$192. Some samples from DAPS are shown in Fig. 5, demonstrating DAPS’s ability to produce diverse samples while preserving the measurement information. We show more samples in Appendix I for further demonstration.
@@
DAPS with large-scale text-conditioned latent diffusion. Due to the flexibility of DAPS, pretrained large-scale text-conditioned latent diffusion models (LDMs) [40, 38, 39] can also serve as the prior model $p({\mathbf{x}}_{0}\mid{\mathbf{c}})$ when given a text prompt ${\mathbf{c}}$. To demonstrate DAPS’s potential, we use Stable Diffusion v2.1 [40] to present LatentDAPS results on three tasks with 768$\times$768 resolution images, as shown in Fig. 1. Additionally, we provide the text prompts used, along with a detailed discussion, in Sec. B.1, where we also include a quantitative evaluation.
@@ 4.3 Further results on MRI
×$\times 4$$\times 8$ PSNR ($\uparrow$) SSIM ($\uparrow$) Meas err ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) Meas err ($\downarrow$) DAPS 31.48 0.762 $1.57$ 29.01 0.681 $1.28$ DPS $26.13$ $0.620$ $9.90$ $20.82$ $0.536$ $6.74$ DiffPIR $28.31$ $0.632$ $10.55$ $26.78$ $0.588$ $7.79$ ScoreMRI [10] $25.97$ $0.468$ $10.83$ $25.01$ $0.405$ $8.36$ CSGM [23] $28.78$ $0.710$ 1.52 $26.15$ $0.625$ 1.14
Table 2: Quantitative results of compressed sensing MRI. We consider ×$\times 4$ and ×$\times 8$ subsampling ratio. Meas err stands for $\ell_{2}$ data consistency error.
@@ 4.4 Ablation Studies
Time efficiency of DAPS. We run DAPS with various configurations to test its performance under different computing budgets. Specifically, we evaluate DAPS with the number of function evaluations (NFE) of the diffusion models ranging from 50 to 4K, with configurations specified in Sec. E.1. Fig. 6(b) shows the relation between the sample quality of DAPS measured by LPIPS with its time cost, which indicates that DAPS is able to generate high-quality samples with a reasonable time cost.
@@
Gaussian deblurring High dynamic range PSNR ($\uparrow$) LPIPS ($\downarrow$) Run time/s PSNR ($\uparrow$) LPIPS ($\downarrow$) Run time/s Gaussian approx. $29.40$ $0.170$ $64.7$ $24.85$ $0.181$ $61.4$ diffusion-score $29.57$ $0.160$ $462.3$ $25.33$ $0.193$ $540.0$ Relative improvement $-0.575\%$ $-6.25\%$ $614.53\%$ $-1.89\%$ $-6.22\%$ $779.48\%$
@@ Appendix A DAPS with Pixel Diffusion Models
Pixel diffusion models learn to approximate the time-dependent score function ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},t)\approx\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$, where $\sigma_{t}$ is a predefined noise schedule with $\sigma_{0}=0$ and $\sigma_{T}=\sigma_{\max}$. Langevin dynamics is used in DAPS due to its simplicity and flexibility. However, other advanced MCMC sampler such as Hamiltonian Monte Carlo (HMC) can be used to enhance the efficiency and quality. We will give detailed derivation in Sec. A.1. We also include a discussion of using the Metropolis Hasting algorithm, a gradient-free MCMC sampling method in Sec. A.2.
@@ B.1 Sampling with Large-scaled Text-conditioned Latent Diffusion Models
[…]
@@
PSLD (SD-v1.5) $0.201$ $30.73$ $0.167$ $0.096$ $30.31$ $0.221$ $30.10$
[…]
Figure 9: Exploring different modes in the posterior distribution by text prompts. Multiple sampling results using LatentDAPS (SD v1.5) were obtained for the super resolution 8× problem on a 512$\times$512 natural image. Text prompts provided additional control over the sampled results..
@@ Appendix C DAPS with Discrete Diffusion Models
[…]
@@ F.1 Inverse Problem Setup
which normalize the data to lies in range $[0,1]$ first. Here $\mathbf{F}$ and $\mathbf{P}$ are discrete Fourier transformation matrices and oversampling matrices with ratio $k/n$. Same as [13], we use an oversampling factor $k=2$ and $n=8$. We normalize input $x_{0}$ by shifting its data range from $[-1,1]$ to $[0,1]$ to better fit practical settings, where the measured signals are usually non-negative.
@@ I.1 More Ablation Study
Recall that we use an ODE sampler to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the estimated mean of the approximated distribution $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$. We use the same number of function evaluations in our ODE sampler throughout the entire algorithm. To test how the number of function evaluations (NFE) in the ODE sampler influences the performance, we try different NFE on two linear tasks and one nonlinear task. In particular, when NFE is $1$, the ODE sampler is equivalent to computing $\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}]$ via Tweedie’s formula. As shown in Fig. 14, increasing NFE in the ODE sampler consistently improves the overall image perceptual quality, but also at the cost of a slightly lower PSNR. This trade-off between PSNR and LPIPS is also observed in [35]. Perceptually, we notice that increasing ODE steps adds more fine-grained details to the produced images, which improves LPIPS but decreases PSNR. This finding is corroborated by Fig. 15, where the reconstructed images appear less blurry and show high-frequency details as the number of ODE steps increases.
@@ I.4 More Qualitative Samples
[…]

### v3

@@ 4.2 Main Results
In addition, our method can generate diversified samples when the measurements contain less information. To illustrate this, we selected two tasks where the posterior distributions may have multiple modes: (1) super-resolution by a factor of 16, and (2) box inpainting with a box size of $\times192$. Some samples from DAPS are shown in Fig. 5, demonstrating DAPS’s ability to produce diverse samples while preserving the measurement information. We show more samples in Appendix I for further demonstration.
@@
DAPS with large-scale text-conditioned latent diffusion. Due to the flexibility of DAPS, pretrained large-scale text-conditioned latent diffusion models (LDMs) [40, 38, 39] can also serve as the prior model $p({\mathbf{x}}_{0}\mid{\mathbf{c}})$ when given a text prompt $\mathbf{c}$}. To demonstrate DAPS’s potential, we use Stable Diffusion v2.1 [40] to present LatentDAPS results on three tasks with $\times768$ resolution images, as shown in Fig. 1. Additionally, we provide the text prompts used, along with a detailed discussion, in Sec. B.1, where we also include a quantitative evaluation.
@@ 4.3 Further results on MRI
$\times 4$ $\times 8$ PSNR ($\uparrow$) SSIM ($\uparrow$) Meas err ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) Meas err ($\downarrow$) DAPS 31.48 0.762 $1.57$ 29.01 0.681 $1.28$ DPS $26.13$ $0.620$ $9.90$ $20.82$ $0.536$ $6.74$ DiffPIR $28.31$ $0.632$ $10.55$ $26.78$ $0.588$ $7.79$ ScoreMRI [10] $25.97$ $0.468$ $10.83$ $25.01$ $0.405$ $8.36$ CSGM [23] $28.78$ $0.710$ 1.52 $26.15$ $0.625$ 1.14
Table 2: Quantitative results of compressed sensing MRI. We consider $\times 4$ and $\times 8$ subsampling ratio. Meas err stands for $\ell_{2}$ data consistency error.
@@ 4.4 Ablation Studies
Time efficiency of DAPS. We run DAPS with various configurations to test its performance under different computing budgets. Specifically, we evaluate DAPS with the number of function evaluations (NFE) of the diffusion models ranging from 50 to 4K, with configurations specified in Sec. E.1. Fig. 6 shows the relation between the sample quality of DAPS measured by LPIPS with its time cost, which indicates that DAPS is able to generate high-quality samples with a reasonable time cost.
@@
Gaussian deblurring High dynamic range PSNR ($\uparrow$) LPIPS ($\downarrow$) Run time/s PSNR ($\uparrow$) LPIPS ($\downarrow$) Run time/s Gaussian approx. $29.40$ $0.170$ $64.7$ $24.85$ $0.181$ $61.4$ diffusion-score $29.57$ $0.160$ $462.3$ $25.33$ $0.193$ $540.0$ relative improvement $0.575\%$ $6.25\%$ $614.53\%$ $1.89\%$ $6.22\%$ $779.48\%$
@@ Appendix A DAPS with Pixel Diffusion Models
Pixel diffusion models learn to approximate the time-dependent score function ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},t)\approx\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$, where $\sigma_{t}$ is a predefined noise schedule with $\sigma_{0}=0$ and $\sigma_{T}=\sigma_{\max}$. Langevin dynamics is used in DAPS due to its simplicity and flexibility. However, other advanced MCMC sampler such as Hamiltonian Monte Carlo (HMC) can be used to enhance the efficiency and quality. We will give a detailed derivation in Sec. A.1. We also include a discussion of using the Metropolis Hasting algorithm, a gradient-free MCMC sampling method in Sec. A.2.
@@ B.1 Sampling with Large-scaled Text-conditioned Latent Diffusion Models
Table 4: Quantitative evaluation on FFHQ $\mathbf{\times}256$ of LatentDAPS with Stable Diffusion v1.5.. The value shows the mean over $100$ images, and all tasks assume the measurement noise level $\beta_{\mathbf{y}}=0.01$. Numbers for PSLD are copied from the original paper.
@@
PSLD (SD-v1.5) $0.201$ $30.73$ $0.167$ - $0.096$ $30.31$ $0.221$ $30.10$ - - - - - - - -
Figure 8: Sampling results of LatentDAPS (SD v1.5) on FFHQ $\mathbf{\times}256$ images. The sampling is enhanced with classifier-free guidance for text with guidance scale $7.5$. The used text prompt is shown in Tab. 5
Figure 9: Exploring different modes in the posterior distribution by text prompts. Multiple sampling results using LatentDAPS (SD v1.5) were obtained for the super resolution 8× problem on a $\times512$ natural image. Text prompts provided additional control over the sampled results.
@@ Appendix C DAPS with Discrete Diffusion Models
PSNR $\uparrow$ Accuracy ($\%$) $\uparrow$ PSNR $\uparrow$ Accuracy ($\%$) $\uparrow$
@@ F.1 Inverse Problem Setup
which normalize the data to lies in range $[0,1]$ first. Here $\mathbf{F}$ and $\mathbf{P}$ are discrete Fourier transformation matrices and oversampling matrices with ratio $k/n$. Same as [13], we use an oversampling factor $k=2$ and $n=8$. We normalize input $x_{0}$ by shifting its data range from [1,1][-1,1] to $[0,1]$ to better fit practical settings, where the measured signals are usually non-negative.
@@ F.3 Baseline Details
DPnP
We direly use the reported time cost and results in DPnP [57].
@@ I.1 More Ablation Study
Recall that we use an ODE sampler to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the estimated mean of the approximated distribution $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$. We use the same number of function evaluations in our ODE sampler throughout the entire algorithm. To test how the number of function evaluations (NFE) in the ODE sampler influences the performance, we try different NFE on two linear tasks and one nonlinear task. In particular, when NFE is 11, the ODE sampler is equivalent to computing $\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}]$ via Tweedie’s formula. As shown in Fig. 14, increasing NFE in the ODE sampler consistently improves the overall image perceptual quality, but also at the cost of a slightly lower PSNR. This trade-off between PSNR and LPIPS is also observed in [35]. Perceptually, we notice that increasing ODE steps adds more fine-grained details to the produced images, which improves LPIPS but decreases PSNR. This finding is corroborated by Fig. 15, where the reconstructed images appear less blurry and show high-frequency details as the number of ODE steps increases.
@@ I.4 More Qualitative Samples
Figure 27: More samples for inpainting of $\mathbf{\times}192$ box. DAPS is able to generate diverse samples when the posterior distribution is multi-modal.